# Smoke and Mirrors in Causal Downstream Tasks

### v3

@@ Abstract
Data: https://doi.org/10.6084/m9.figshare.26484934.v2
@@ 1 Introduction
  • We theoretically show how many design choices can affect the answer to a causal question, from the data used for training, the architecture choices, and even seemingly innocuous standard practices like thresholding the predictions into hard labels, or using held out accuracy for model selection (a common practice in many AI for science benchmarks, e.g., (Sun et al., 2023)). To facilitate future research on representation learning for causal downstream tasks, we formulate the representation desiderata to obtain accurate estimates for downstream causal queries together with best practices.
  • To showcase the practical impact of these design choices, we design and collect a new dataset, ISTAnt, from a real randomized controlled trial, reflecting a real-world pipeline in experimental behavioral ecology, which we will release to accelerate research on representation learning for causal downstream tasks. To the best of our knowledge, this is the first real-world data set specifically designed for causal inference downstream tasks from high-dimensional observations.
  • On our dataset, we fine-tune 6 480 state-of-the-art methods (Dosovitskiy et al., 2020; Zhai et al., 2023; Radford et al., 2021; He et al., 2022; Oquab et al., 2023) in the few- and many-shot settings. Empirically, we confirm that the seemingly innocuous design choices like which samples to annotate, which model to use, whether or not to threshold the labels, and how to do model selection have a major impact on the accuracy of the causal estimate. Since our ground-truth estimate of the causal effect depends on the trial’s design, we propose a new synthetic benchmark based on MNIST (LeCun, 1998) controlling for the causal model, and we replicated the analysis.
@@ Acknowledgments
We thank Piersilvio De Bartolomeis, and the full Causal Learning and Artificial Intelligence (CLAI) group at ISTA for the extremely helpful discussions. Riccardo Cadei was supported by a Google Research Scholar Award and a Google Initiated Gift to Francesco Locatello. We thank the Social Immunity team at ISTA, particularly Michaela Hönigsberger and Wilfrid Jean Louis, for supporting the ecological experiment and Farnaz Beikzadeh Abbasi, Luisa Fiebig and Martin Estermann for annotating ant behavior in ISTAnt.
@@ Appendix C ISTAnt
In our study, we analyzed grooming behavior in the ant Lasius neglectus in groups of three worker ants. The workers for the experiment were obtained from their laboratory stock colony, which had been collected from the field in 2022 in the Botanical Garden Jena, Germany. Ant collection and all experimental work were performed in compliance with international, national and institutional regulations and ethical guidelines. For the experiment, the body surface of one of the three ants was treated with a suspension of either of two microparticle types (diameter  5 µm) to induce grooming by the two nestmates, which were individually color-coded by application of a dot of blue or orange paint, respectively. The three ants were housed in small plastic containers (diameter 28mm, height 30mm) with moistened, plastered ground and the interior walls covered with PTFE (polytetrafluoroethane) to hamper climbing by the ants. Filming occurred in a temperature- and humidity-controlled room at 23°C within a custom-made filming box with controlled lighting and ventilation conditions. We set up nine ant groups at a time (always containing both treatments) and placed them randomly on positions 1-9 marked on the floor in a 3x3 grid with a distance of about 3mm from each other. Figure 7 illustrates the filming box and the displaying of the containers in each batch. The experiment was performed on two consecutive days. Videos were acquired using a USB camera (FLIR blackfly S BFS-U3-120S4C, Teledyne FLIR) with a high-performance lens (HP Series 25mm Focal Length, Edmund optics 86-572) in OBS studio 29.0.0 [Bailey, 2017] at a framerate of 30 FPS and a resolution of 2500x2500 pixels. From each original video (105x105 mm), we generated 9 individual videos .mkv (each  32x32 mm, 770x770 pixels) by determining exact coordinates per container from one frame in GIMP 2.10.36 [Kimball and Mattis, 2023] and cropping of the videos with FFmpeg 6.1.1 [Tomar, 2006]. Annotation was performed over two consecutive days by three observers who had not been involved in the experimental setup or recording and were unaware of the treatment assignments to ensure bias-free behavioral annotation. They annotated the behavior of the ants during video observations, using custom-made software that saves the start and end frames of behaviors marked in a .csv file. In one of the videos, one of the nestmates’ legs got inadvertently stuck to its body surface during the color-coding, interfering with its behavior, so the video was discarded. This left 44 videos from 5 independent setups (n=24 of treatment 1 and n=20 of treatment 2) of 10 minutes each for a total of 792 000 annotated frames. For each video, we provide the following information: the number of the set to which it belongs (1-5); the number of the position within the set reflecting the position of the ant group under the camera (1-9), for which we also provide ‘coordinates’ in the 3x3 grid (taking values -1/0/1 for both X and Y axis); treatment (1 or 2); the hour of the day when the recording was started (in 24h CEST); experimental day (A or B); the top left coordinate of the cropping square from the original video (CropX/CropY); the person annotating the video (given as A, B, C); the date of annotation (1: first day, 2: second day) and in which order the videos were annotated by each person (both reflecting a possible training effect of the person).
((a)) Camera set-up
((b)) Filming box
((c)) Containers grid

### v4

@@ Abstract
Data: https://doi.org/10.6084/m9.figshare.26484934.v3
@@ 1 Introduction
  • We theoretically show how many design choices can affect the answer to a causal question, from the data used for training, the architecture choices, and even seemingly innocuous standard practices like thresholding the predictions into hard labels, or using held-out accuracy for model selection (a common practice in many AI for science benchmarks, e.g., (Sun et al., 2023)). To facilitate future research on representation learning for causal downstream tasks, we formulate the representation desiderata to obtain accurate estimates for downstream causal queries together with best practices.
  • To showcase the practical impact of these design choices, we design and collect a new dataset, ISTAnt, from a real randomized controlled trial, reflecting a real-world pipeline in experimental behavioral ecology, which we release to accelerate research on representation learning for causal downstream tasks. To the best of our knowledge, this is the first real-world data set specifically designed for causal inference downstream tasks from high-dimensional observations.
  • On our dataset, we fine-tune 6 480 state-of-the-art models (Dosovitskiy et al., 2020; Zhai et al., 2023; Radford et al., 2021; He et al., 2022; Oquab et al., 2023) in the few- and many-shot settings. Empirically, we confirm that the seemingly innocuous design choices like which samples to annotate, which model to use, whether or not to threshold the labels, and how to do model selection have a major impact on the accuracy of the causal estimate. Since our ground-truth estimate of the causal effect depends on the trial’s design, we propose a new synthetic benchmark based on MNIST (LeCun, 1998) controlling for the causal model, and we replicated the analysis.
@@ Acknowledgments
We thank Piersilvio De Bartolomeis, and the full Causal Learning and Artificial Intelligence (CLAI) group at ISTA for the extremely helpful discussions. We thank the Social Immunity team at ISTA, particularly Michaela Hönigsberger and Wilfrid Jean Louis, for supporting the ecological experiment, Jinook Oh for the annotation program and Farnaz Beikzadeh Abbasi, Luisa Fiebig, and Martin Estermann for annotating ant behavior in ISTAnt. Riccardo Cadei was supported by a Google Research Scholar Award and a Google Initiated Gift to Francesco Locatello.
@@ Appendix C ISTAnt
In our study, we analyzed grooming behavior in the ant Lasius neglectus in groups of three worker ants. The workers for the experiment were obtained from their laboratory stock colony, which had been collected from the field in 2022 in the Botanical Garden Jena, Germany. Ant collection and all experimental work were performed in compliance with international, national and institutional regulations and ethical guidelines. For the experiment, the body surface of one of the three ants was treated with a suspension of either of two microparticle types (diameter  5 µm) to induce grooming by the two nestmates, which were individually color-coded by application of a dot of blue or orange paint, respectively. The three ants were housed in small plastic containers (diameter 28mm, height 30mm) with moistened, plastered ground and the interior walls covered with PTFE (polytetrafluoroethane) to hamper climbing by the ants. Filming occurred in a temperature- and humidity-controlled room at 23°C within a custom-made filming box with controlled lighting and ventilation conditions. We set up batches of nine ant groups at a time (always containing both treatments) and placed them randomly on positions 1-9 marked on the floor in a 3x3 grid, about 3mm from each other. Figure 7 illustrates the filming box and the displaying of the containers in each batch. The experiment was performed on two consecutive days. Videos were acquired using a USB camera (FLIR blackfly S BFS-U3-120S4C, Teledyne FLIR) with a high-performance lens (HP Series 25mm Focal Length, Edmund optics 86-572) in OBS studio 29.0.0 [Bailey, 2017] at a framerate of 30 FPS and a resolution of 2500x2500 pixels. From each original video (105x105 mm), we generated nine individual videos .mkv (each  32x32 mm, 770x770 pixels) by determining exact coordinates per container from one frame in GIMP 2.10.36 [Kimball and Mattis, 2023] and cropping of the videos with FFmpeg 6.1.1 [Tomar, 2006]. Annotation was performed over two consecutive days by three observers who had not been involved in the experimental setup or recording and were unaware of the treatment assignments to ensure bias-free behavioral annotation. They annotated the ants’ behavior by video observations, using custom-made software that saves the start and end frames of behaviors marked in a .csv file (see ’annotations’ folder). In one of the videos, one of the nestmates’ legs got inadvertently stuck to its body surface during the color-coding, interfering with its behavior, so the video was discarded. This left 44 videos from 5 independent setups (n=24 of treatment 1 and n=20 of treatment 2) of 10 minutes each for a total of 792 000 annotated frames. For each video, we provide the following information: the batch number (1-5); the position number within the batch (1-9), for which we also provide the ‘coordinates’ in the 3x3 grid (taking values -1/0/1 for both X and Y axis); treatment (1 or 2); the hour of the day when the recording was started (in 24h CEST); experimental day (A or B); the person annotating the video (given as A, B, C) and in which order the videos were annotated by each person, capturing a potential training effect of the person.
((a)) Filming box
((b)) Batch example